OpenStudy (jennyrlz):

help ; ; Find the vertex, focus, directrix, and focal width of the parabola. -1/40x^2=y

OpenStudy (jennyrlz):

@geerky42 i cant find the focal width

OpenStudy (misty1212):

HI!!

OpenStudy (jennyrlz):

hello xD

geerky42 (geerky42):

Not very familiar with it, sorry.

OpenStudy (jennyrlz):

i know the vertex is 0.0

OpenStudy (misty1212):

focal width of \(4py=x^2\) is \(|4p|\) that is all, you don't really find it, it is right there

OpenStudy (jennyrlz):

um not in my answer choices, ive narrowed it down to two

OpenStudy (jennyrlz):

either 40 or 160

OpenStudy (misty1212):

\[-\frac{1}{40}x^2=y\\ -40y=x^2\] is a nice way to write the parabola, you get the vertex directrix, focus, and focal width by looking

OpenStudy (misty1212):

it is \(40\) not \(4\times 40\)

OpenStudy (jennyrlz):

so its 160 xD

OpenStudy (misty1212):

NO!!

OpenStudy (misty1212):

just \(40\)

OpenStudy (jennyrlz):

oh i miss interpreted you answer xD

OpenStudy (jennyrlz):

i got 4x40 lol

OpenStudy (misty1212):

you get the vertex as \((0,0)\) right?

OpenStudy (jennyrlz):

yup

OpenStudy (misty1212):

how about the rest ?

OpenStudy (jennyrlz):

and focus as 0,-10

OpenStudy (jennyrlz):

and the directrix of y as 10

OpenStudy (misty1212):

yup

OpenStudy (misty1212):

yup

OpenStudy (jennyrlz):

i might need your help again so be attentive xD

OpenStudy (misty1212):

look like you got it all